# Rise of the forsaken relics: connecting present-day stellar streams and phase-mixed galaxies to the Epoch of Reionization

### v1

@@ Abstract
1 Introduction
The Epoch of Reionization (EoR) is the time period when the first stars, galaxies, and accreting black holes emitted ultraviolet (UV) radiation and gradually photo-ionized neutral hydrogen in the intergalactic medium (IGM) throughout the universe (Barkana & Loeb, 2001; Furlanetto et al., 2006). The timing and nature of this process remain highly uncertain, while the properties of the ionizing sources are currently unknown. The most likely scenario is that the UV radiation from star-forming galaxies drives reionization, with the process completing around $z\sim 6$ (Shapiro et al., 1994; Madau et al., 1999; Faucher-Giguère et al., 2008; Becker & Bolton, 2013), although precisely which galaxies dominate is unclear. The Hubble Space Telescope (HST) has directly observed faint reionization-era galaxies down to $M_{\mathrm{UV}}\sim-13$ in fields behind lensing clusters (Finkelstein et al., 2015; Bouwens et al., 2017; Livermore et al., 2017; Atek et al., 2018; Bouwens et al., 2022b), albeit with sizable uncertainties at the faint end. Ongoing and future observations with the James Webb Space Telescope (JWST) are pushing these measurements to $z\gtrsim 10$ and slightly fainter magnitudes (Finkelstein et al., 2023; Leung et al., 2023; Navarro-Carrera et al., 2023; Pérez-González et al., 2023; Atek et al., 2024; Finkelstein et al., 2024; Adams et al., 2024; Harikane et al., 2025; Chemerynska et al., 2025).
However, some reionization models predict that fainter yet more numerous low-mass galaxies, with, e.g. $M_{\mathrm{UV}}\sim-15$ to 8-8, are major contributors to the total budget of ionizing photons (Kuhlen & Faucher-Giguère, 2012; Wise et al., 2014; Robertson et al., 2015; Atek et al., 2015; Ma et al., 2018a, b; Finkelstein et al., 2019; Atek et al., 2024; Mascia et al., 2024; Simmonds et al., 2024; Harshan et al., 2024; Wu & Kravtsov, 2024). Many of these sources will remain beyond the reach of even JWST (Boylan-Kolchin et al., 2015, 2016; Weisz & Boylan-Kolchin, 2017). While observations of lensed galaxies can probe fainter magnitudes (down to $M_{\mathrm{UV}}\sim-12$), they introduce significant systematic uncertainties due to magnification effects (Atek et al., 2018; Bouwens et al., 2022a). Additionally, low-mass galaxies provide valuable insights into the impact of stellar feedback and the nature of dark matter on galaxy formation, particularly at high redshifts (Sun & Furlanetto, 2016; Bullock & Boylan-Kolchin, 2017; Sales et al., 2022; Sipple & Lidz, 2024; Sipple et al., 2025). Novel approaches for probing the faint/low-mass-end of the $z\gtrsim 6$ galaxy UV luminosity function (UVLF)/stellar mass function (SMF) are hence valuable - not only in determining the role of low-luminosity galaxies in reionizing the universe, but also in understanding feedback processes and the properties of dark matter.
@@ 1 Introduction
In this paper, we investigate these issues using a suite of simulated MW/LG-like galaxies from zoomed cosmological hydrodynamic simulations. In section 2, we describe the simulations used in this work. Section 3 discusses our criteria for classifying present-day substructures and our methodology for tracking them back to their simulated high-$z$ progenitors. Through Sections 4 to 8, we illustrate the impact of adding disrupted galaxies to the near-far technique and also test the feasibility of using them by imposing detectability constrains. Finally, we summarize these results in Section 9.
2 Simulations
2.1 FIRE-2 simulations
We use cosmological zoom-in baryonic simulations from the Feedback in Realistic Environments (FIRE-2) project (Wetzel et al., 2023, 2025)  [^1] https://fire.northwestern.edu/. They are run with the Gizmo code (Hopkins, 2015; Hopkins et al., 2018a), which uses a mesh-free, finite volume Lagrangian method for accurate hydrodynamics. Baryonic physical processes like star formation and stellar feedback are implemented using the FIRE-2 numerical prescriptions (Hopkins et al., 2018a). The multiphase interstellar medium (ISM) in each simulated galaxy is captured with realistic radiative heating and cooling processes across temperatures from 10 K to 1010 K.
Simulated star particles form out of gas parcels satisfying a Jeans instability criterion, and inherit the mass and elemental abundances from their respective gas cells. Each star particle in FIRE represents a single-age, single-abundance stellar population, whose evolution is based on stellar population synthesis models from STARBURST99 (Leitherer et al., 1999). FIRE also models stellar feedback processes like core-collapse and Typa Ia supernovae, radiation pressure, continuous mass loss, photoionization and photo-electric heating in a time-resolved manner. Furthermore, FIRE simulations implement a spatially uniform UV/X-ray ionizing background that evolves with redshift (Faucher-Giguère et al., 2009). Some of the caveats of this model are that it neglects the inhomogeneous nature of reionization and also has an earlier reionization time ($z\sim 10$) than recent measurements suggest ($z\sim 7$) (Planck Collaboration et al., 2020). But all the simulations that we use in this work self-consistently employ this model and hence will not result in any simulation-to-simulation difference due to these caveats. Additionally, these simulations includes spurious heating of neutral gas at $\lesssim$1000 K at $z\gtrsim 10$ due to cosmic rays, which suppresses star formation in low-mass halos at high redshifts. But this has no effect after reionization begins and hence our results at $z\lesssim 9$ are self-consistent (Gandhi et al., 2024). For a more detailed description of the physics implemented in FIRE simulations, see Hopkins et al. (2018a, b).
@@ 2.1 FIRE-2 simulations
2.2 Identifying low-mass galaxies
The DM halos and subhalos (smaller virialized halos within a larger halo) in each snapshot of the core simulations are identified using the ROCKSTAR 6D phase-space halo finder (Behroozi et al., 2013a) to create a halo catalog for each of the 600 snapshots. The (sub)halos are identified according to their radius, such that it encloses 200 times the mean matter density of the Universe ($R_{\rm 200m}$), and only the ones with bound mass fraction >> 0.4 and at least 30 DM particles are kept. This criterion reduces low-resolution artifacts, particularly contamination from the low-resolution DM particles (the low-resolution particles are a consequence of the zoom-in re-simulation technique employed). After the halo catalogs have been formed, a merger tree is constructed by connecting the (sub)halos in time using consistent-trees (Behroozi et al., 2013b). To ensure numerical stability, the halo catalogs and merger trees are constructed using only DM particles (Samuel et al., 2020).
@@ 2.2 Identifying low-mass galaxies
This method is conservative enough to keep galaxy properties stable across time at lower redshifts, and excludes nearly all stream stars (Panithanpaisal et al., 2021; Samuel et al., 2020). These assignments are then used as a starting point for identifying streams, as described in Section 3.1. However, these criteria can be too stringent when assigning star particles to low-mass (sub)halos at $z\geq 6$, which in turn can lead to an incomplete low-mass (sub)halo population. In order to improve the overall completeness of the star particle assignment to each (sub)halo at $z\geq 6$, we follow the method implemented in Gandhi et al. (2024): we change the enclosing radius from $0.8\,R_{\rm 200m}$ to $1.0\,R_{\rm 200m}$ and do not apply any constraints on the velocity relative to $V_{\rm circ,max}$. Furthermore, we keep all the halos with at least two star particles instead of six, since most of the galaxies that we analyze at high redshifts are progenitors of well-resolved galaxies at $z=0$.
@@
Note. — Simulation name: References listed below; top 7 are isolated MW-like and bottom three are paired MW/M31-like. $M_{\rm 200m}$ and $R_{\rm 200m}$: total mass and spherical radius within which the mean density is 200 times the matter density of the universe. $M_{\star,90}$: stellar mass within a spherical radius that encloses 90% of the stellar mass within 20 kpc. $z_{\mathrm{MR_{3:1}}}$: the redshift at which the stellar mass ratio of the most massive halo to the second most massive subhalo is 3:1, which is when the dominant host galaxy emerges. Numbers of surviving satellite galaxies ($M_{\star}\gtrsim 10^{4.5}M_{\odot}$), stellar streams ($M_{\star}\gtrsim 10^{6}M_{\odot}$), phase-mixed galaxies ($M_{\star}\gtrsim 10^{6}M_{\odot}$), and field galaxies ($M_{\star}\gtrsim 10^{4.5}M_{\odot}$) are determined as described in Section 3.1.
@@
3 Present-day substructures and their high redshift progenitors
3.1 Substructure classification algorithm
@@
Figure 1: Present-day substructures classified by our algorithm, with manual reclassification. The center of the host galaxy is denoted by the black cross in each panel. Top: Three examples correctly classified by the automated algorithm described in Section 3.1. Left: a satellite within $R_{\rm 200m}$ (black dashed circle) and a field galaxy outside $R_{\rm 200m}$ Center: a stellar stream. Right: a phase-mixed galaxy. Bottom: Manual reclassification of substructures with complex morphologies. Left: a satellite galaxy which was classified as a stream by the algorithm due to a handful of unbound star particles. Center: a stream which was classified as a phase-mixed galaxy due to its large mass. Right: a phase-mixed galaxy which was classified as a satellite galaxy due to its small size.
@@
We use three criteria to select and classify present-day substructures. We first require that the number of star particles be larger than 100 ($M_{*}\gtrsim 5\times 10^{5}\,M_{\odot}$). [^3] We manually reduce the stellar mass requirement to $10^{4.5}$ $M_{\odot}$ for the case of satellite galaxies to be consistent with Gandhi et al. (2024). This choice seems reasonable since these galaxies are more compact than streams and therefore easier to detect. In Section 8 we quantify detectability in terms of surface brightness. To be considered disrupted rather than intact, the maximum pairwise separation between any two star particles must also be greater than 120 kpc. Finally, the local velocity dispersion $\sigma$ of each disrupted substructure is compared to a stellar-mass dependent threshold $\sigma_{\mathrm{th}}(M_{*})$ (Panithanpaisal et al., 2021, Eq. 2). Substructures with $\sigma<\sigma_{\mathrm{th}}(M_{*})$ are considered streams, while those with $\sigma>\sigma_{\mathrm{th}}(M_{*})$ are considered phase-mixed galaxies. In some borderline cases where a substructure barely satisfies some criterion, but looks distinctively like another category, we reclassify the substructure manually. Table 1 shows the number of different substructures in each of the FIRE-2 simulations, along with the properties of the host(s). We show examples of the substructures classified by the algorithm as satellite galaxies, stellar streams, phase-mixed galaxies, and field galaxies, in Figure 1, along with cases where we manually reclassify the original assignment according to the algorithm.
@@
3.2 Identifying the high-redshift progenitors
@@
4 Stellar mass distribution of progenitor galaxies at high redshifts
@@
4.1 Stellar mass budget during the EoR
@@
In the isolated suite, the progenitors of phase-mixed galaxies contribute the most to the total stellar mass budget, followed by the stellar stream progenitors. In the paired suite, progenitors of phase-mixed galaxies again contribute the most to the total stellar mass budget, followed by the progenitors of field galaxies, streams, and satellites, respectively. Though the stream progenitors contribute less than that of the field galaxies in the paired suite, in both cases, the progenitors of the present-day disrupted galaxies yield more stellar mass during the EoR than the progenitors of intact low-mass galaxies. Even though the stream progenitors contribute less than phase-mixed galaxies, they play a larger role than intact low-mass galaxies. The fractional stellar mass contribution from progenitors of each classification, relative to progenitors of the main galaxy, is roughly constant from z=69z=6-9, with disrupted galaxies contributing at least as much as intact satellites across this entire range (Figure 2, bottom panels). There is no strong mass-dependent bias across different present-day substructures. This suggests that neglecting the progenitors of streams and phase-mixed galaxies can underestimate the reionization-era ionizing photon budget derived from the near-far technique. However, even with adding the disrupted galaxies we can recover only up to $\sim$50% of the total budget.
@@
4.2 Hierarchical structure formation at high redshifts
@@
5 Stellar Mass Functions and UV Luminosity Functions at High-Redshifts
@@
These procedures result in reconstructed high-$z$ SMFs/UVLFs for both the all progenitors and fossil record scenarios. We compare them to universal benchmark SMFs/UVLFs in sections 5.2 and 5.3.
Figure 4: Present-day disrupted and intact galaxies recover different differential SMFs and UVLFs at $z=7$. All panels show median reconstructed SMFs (top) and UVLFs (bottom) for the progenitors of various present-day substructure types, without (solid curves) and with (dashed curves) detectability constraints (Section 8). The isolated galaxy suite is shown in the left column; paired galaxies in the right column. The proto-MW/LG (black), which includes the progenitors of the present-day main host(s), is compared with reconstructions from the progenitors of satellite galaxies (red), stellar streams (blue), phase-mixed galaxies (green), and field galaxies (yellow).
5.1 Reconstructed differential SMFs and UVLFs for each present-day classification
Figure 4 compares the differential SMFs (top panel) and UVLFs (bottom panel) of the proto-MW/LG (black) at $z=7$ with the ones constructed from the progenitors of present-day satellite galaxies (red), stellar streams (blue), phase-mixed galaxies (green), and field galaxies (yellow). The solid lines show reconstructions from the full set of progenitors and the dashed lines show reconstructions obtained using a simple model for the detectability of present-day substructures with the LSST (section 8). We show results for the isolated suite and the paired suite in the left and right panels, respectively. The comparison between the two panels illustrates how effective the near-far technique is when applied to each of the MW and the LG alone. The curves show the median across all of the simulations in the respective galaxy suite. To highlight differences between classifications, the SMFs/UVLFs are computed in the all progenitors scenario rather than the fossil record scenario.

### v2

@@ Abstract
I Introduction
The Epoch of Reionization (EoR) is the time period when the first stars, galaxies, and accreting black holes emitted ultraviolet (UV) radiation and gradually photo-ionized neutral hydrogen in the intergalactic medium (IGM) throughout the universe (Barkana & Loeb, 2001; Furlanetto et al., 2006). The timing and nature of this process remain highly uncertain, while the properties of the ionizing sources are currently unknown. The most likely scenario is that the UV radiation from star-forming galaxies drives reionization, with the process completing around $z\sim 6$ (Shapiro et al., 1994; Madau et al., 1999; Faucher-Giguère et al., 2008; Becker & Bolton, 2013), although precisely which galaxies dominate is unclear. The Hubble Space Telescope (HST) has directly observed faint reionization-era galaxies down to $M_{\mathrm{UV}}\sim-13$ in fields behind lensing clusters (Finkelstein et al., 2015; Bouwens et al., 2017; Livermore et al., 2017; Atek et al., 2018; Bouwens et al., 2022b), albeit with sizable uncertainties at the faint end. Ongoing and future observations with the James Webb Space Telescope (JWST) are pushing these measurements to $z\gtrsim 10$ and slightly fainter magnitudes (e.g., $M_{\mathrm{UV}}\sim-12$) (Finkelstein et al., 2023; Leung et al., 2023; Navarro-Carrera et al., 2023; Pérez-González et al., 2023; Atek et al., 2024; Finkelstein et al., 2024; Adams et al., 2024; Harikane et al., 2025; Chemerynska et al., 2025; Atek et al., 2025).
However, some reionization models predict that fainter yet more numerous low-mass galaxies are major contributors to the total budget of ionizing photons, potentially with $M_{\mathrm{UV}}$ down to $\sim-8$ (Kuhlen & Faucher-Giguère, 2012; Wise et al., 2014; Robertson et al., 2015; Atek et al., 2015; Ma et al., 2018a, b; Finkelstein et al., 2019; Atek et al., 2024; Mascia et al., 2024; Simmonds et al., 2024; Harshan et al., 2024; Wu & Kravtsov, 2024). Many of these sources will remain beyond the reach of even JWST (Boylan-Kolchin et al., 2015, 2016; Weisz & Boylan-Kolchin, 2017). While observations of lensed galaxies can probe fainter magnitudes (down to $M_{\mathrm{UV}}\sim-12$), they introduce significant systematic uncertainties due to magnification effects (Atek et al., 2018; Bouwens et al., 2022a). Additionally, low-mass galaxies provide valuable insights into the impact of stellar feedback and the nature of dark matter on galaxy formation, particularly at high redshifts (Sun & Furlanetto, 2016; Bullock & Boylan-Kolchin, 2017; Sales et al., 2022; Sipple & Lidz, 2024; Sipple et al., 2025). Novel approaches for probing the faint/low-mass-end of the $z\gtrsim 6$ galaxy UV luminosity function (UVLF)/stellar mass function (SMF) are hence valuable not only in determining the role of low-luminosity galaxies in reionizing the universe, but also in understanding feedback processes and the properties of dark matter.
@@ I Introduction
In this paper, we investigate these issues using a suite of simulated MW/LG-like galaxies from zoomed cosmological hydrodynamic simulations. In section II, we describe the simulations used in this work. Section III discusses our criteria for classifying present-day substructures and our methodology for tracking them back to their simulated high-$z$ progenitors. Through Sections IV to VIII, we illustrate the impact of adding disrupted galaxies to the near-far technique and also test the feasibility of using them by imposing detectability constrains. Finally, we summarize these results in Section IX.
II Simulations
II.1 FIRE-2 simulations
We use cosmological zoom-in baryonic simulations from the Feedback in Realistic Environments (FIRE-2) project (Wetzel et al., 2023, 2025)  [^1] https://fire.northwestern.edu/. They are run with the Gizmo code (Hopkins, 2015; Hopkins et al., 2018a), which uses a mesh-free, finite mass (MFM) Lagrangian method for accurate hydrodynamics. Baryonic physical processes like star formation and stellar feedback are implemented using the FIRE-2 numerical prescriptions (Hopkins et al., 2018a). The multiphase interstellar medium (ISM) in each simulated galaxy is captured with realistic radiative heating and cooling processes across temperatures from 10 K to 1010 K.
Simulated star particles form out of gas parcels satisfying a Jeans instability criterion, and inherit the mass and elemental abundances from their respective gas cells. Each star particle in FIRE represents a single-age, single-abundance stellar population, whose evolution is based on stellar population synthesis models from STARBURST99 (Leitherer et al., 1999) assuming a Kroupa initial mass function (IMF) (Kroupa, 2001). FIRE also models stellar feedback processes like core-collapse and Typa Ia supernovae, radiation pressure, continuous mass loss, photoionization and photo-electric heating in a time-resolved manner. Furthermore, FIRE simulations implement a spatially uniform UV/X-ray ionizing background that evolves with redshift (Faucher-Giguère et al., 2009). Some of the caveats of this model are that it neglects the inhomogeneous nature of reionization and also has an earlier reionization time ($z\sim 10$) than recent measurements suggest ($z\sim 7$) (Planck Collaboration et al., 2020). But all the simulations that we use in this work self-consistently employ this model and hence will not result in any simulation-to-simulation difference due to these caveats. Additionally, these simulations includes spurious heating of neutral gas at $\lesssim$1000 K at $z\gtrsim 10$ due to cosmic rays, which suppresses star formation in low-mass halos at high redshifts. But this has no effect after reionization begins and hence our results at $z\lesssim 9$ are self-consistent (Gandhi et al., 2024). For a more detailed description of the physics implemented in FIRE simulations, see Hopkins et al. (2018a, b).
@@ II.1 FIRE-2 simulations
II.2 Identifying low-mass galaxies
The DM halos and subhalos (smaller virialized halos within a larger halo) in each snapshot of the core simulations are identified using the ROCKSTAR 6D phase-space halo finder (Behroozi et al., 2013a) to create a halo catalog for each of the 600 snapshots. The (sub)halos are identified according to their radius, such that it encloses 200 times the mean matter density of the Universe ($R_{\rm 200m}$), and only the ones with bound mass fraction >> 0.4 and at least 30 DM particles are kept. The bound mass fraction serves to select friends-of-friends groups in phase space that have sufficient density to be gravitationally self-bound. We take a value slightly lower than ROCKSTAR’s default of 0.5 since we are interested in finding the nuclei of partially disrupted halos for further tracking, but well above the threshold of 0.15 where the halo mass function is sensitive to changes in this parameter (Behroozi et al., 2013a). After the halo catalogs have been formed, a merger tree is constructed by connecting the (sub)halos in time using consistent-trees (Behroozi et al., 2013b). To ensure numerical stability, the halo catalogs and merger trees are constructed using only DM particles (Samuel et al., 2020).
@@ II.2 Identifying low-mass galaxies
This method is conservative enough to keep galaxy properties stable across time at lower redshifts, and excludes nearly all stream stars (Panithanpaisal et al., 2021; Samuel et al., 2020). These assignments are then used as a starting point for identifying streams, as described in Section III.1. However, these criteria can be too stringent when assigning star particles to low-mass (sub)halos at $z\geq 6$, which in turn can lead to an incomplete low-mass (sub)halo population. In order to improve the overall completeness of the star particle assignment to each (sub)halo at $z\geq 6$, we follow the method implemented in Gandhi et al. (2024): we change the enclosing radius from $0.8\,R_{\rm 200m}$ to $1.0\,R_{\rm 200m}$ and do not apply any constraints on the velocity relative to $V_{\rm circ,max}$. Furthermore, we keep all the halos with at least two star particles instead of six, since most of the galaxies that we analyze at high redshifts are progenitors of well-resolved galaxies at $z=0$.
@@
Note. — Simulation name: References listed below; top 7 are isolated MW-like and bottom three are paired MW/M31-like. $M_{\rm 200m}$ and $R_{\rm 200m}$: total mass and spherical radius within which the mean density is 200 times the matter density of the universe. $M_{\star,90}$: stellar mass within a spherical radius that encloses 90% of the stellar mass within 20 kpc. $z_{\mathrm{MR_{3:1}}}$: the redshift at which the stellar mass ratio of the most massive halo to the second most massive subhalo is 3:1, which is when the dominant host galaxy emerges. Numbers of surviving satellite galaxies ($M_{\star}\gtrsim 10^{4.5}M_{\odot}$), stellar streams ($M_{\star}\gtrsim 10^{5.5}M_{\odot}$), phase-mixed galaxies ($M_{\star}\gtrsim 10^{5.5}M_{\odot}$), and field galaxies ($M_{\star}\gtrsim 10^{4.5}M_{\odot}$) are determined as described in Section III.1.
@@
III Present-day substructures and their high redshift progenitors
III.1 Substructure classification algorithm
@@
Figure 1: Present-day substructures classified by our algorithm, with manual reclassification. The center of the host galaxy is denoted by the black cross in each panel. Top: Three examples correctly classified by the automated algorithm described in Section III.1. Left: a satellite within $R_{\rm 200m}$ (black dashed circle) and a field galaxy outside $R_{\rm 200m}$ Center: a stellar stream. Right: a phase-mixed galaxy. Bottom: Manual reclassification of substructures with complex morphologies. Left: a satellite galaxy which was classified as a stream by the algorithm due to a handful of unbound star particles. Center: a stream which was classified as a phase-mixed galaxy due to its large mass. Right: a phase-mixed galaxy which was classified as a satellite galaxy due to its small size.
@@
We use three criteria to select and classify present-day substructures. We first require that the number of star particles be larger than 100 ($M_{*}\gtrsim 5\times 10^{5}\,M_{\odot}$). [^3] We manually reduce the stellar mass requirement to $10^{4.5}$ $M_{\odot}$ for the case of satellite galaxies to be consistent with Gandhi et al. (2024). This choice seems reasonable since these galaxies are more compact than streams and therefore easier to resolve and detect. In Section VIII we quantify detectability in terms of surface brightness. This ensures that the streams have enough star particles to be resolved in our simulations. To be considered disrupted rather than intact, the maximum pairwise separation between any two star particles must also be greater than 120 kpc. Finally, the local velocity dispersion $\sigma$ of each disrupted substructure is compared to a stellar-mass dependent threshold $\sigma_{\mathrm{th}}(M_{*})$ (Panithanpaisal et al., 2021, Eq. 2). Substructures with $\sigma<\sigma_{\mathrm{th}}(M_{*})$ are considered streams, while those with $\sigma>\sigma_{\mathrm{th}}(M_{*})$ are considered phase-mixed galaxies. In some borderline cases where a substructure barely satisfies some criterion, but looks distinctively like another category, we reclassify the substructure manually. Table 1 shows the number of different substructures in each of the FIRE-2 simulations, along with the properties of the host(s). We show examples of the substructures classified by the algorithm as satellite galaxies, stellar streams, phase-mixed galaxies, and field galaxies, in Figure 1, along with cases where we manually reclassify the original assignment according to the algorithm.
@@
III.2 Identifying the high-redshift progenitors
@@
IV Stellar mass distribution of progenitor galaxies at high redshifts
@@
IV.1 Stellar mass budget during the EoR
@@
In the isolated suite, the progenitors of phase-mixed galaxies contribute the most to the total stellar mass budget, followed by the stellar stream progenitors. In the paired suite, progenitors of phase-mixed galaxies again contribute the most to the total stellar mass budget, followed by the progenitors of field galaxies, streams, and satellites, respectively. Though the stream progenitors contribute less than that of the field galaxies in the paired suite, in both cases, the progenitors of the present-day disrupted galaxies yield more stellar mass during the EoR than the progenitors of intact low-mass galaxies. Even though the stream progenitors contribute less than phase-mixed galaxies, they play a larger role than intact low-mass galaxies. The fractional stellar mass contribution from progenitors of each classification, relative to progenitors of the main galaxy, is roughly constant from z=69z=6-9, with disrupted galaxies contributing at least as much as intact satellites across this entire range (Figure 2, bottom panels). There is no strong mass-dependent bias across different present-day substructures. This suggests that neglecting the progenitors of streams and phase-mixed galaxies can underestimate the reionization-era ionizing photon budget derived from the near-far technique. However, even with adding the disrupted galaxies we can recover only up to $\sim$50% of the total budget. The rest of the stellar mass is encompassed by the progenitors of the present-day main host galaxy (Horta et al., 2024).
@@
IV.2 Hierarchical structure formation at high redshifts
@@
V Stellar Mass Functions and UV Luminosity Functions at High-Redshifts
@@
These procedures result in reconstructed high-$z$ SMFs/UVLFs for both the all progenitors and fossil record scenarios. We compare them to universal benchmark SMFs/UVLFs in sections V.2 and V.3.
Figure 4: Present-day disrupted and intact galaxies recover different differential SMFs and UVLFs at $z=7$. All panels show median reconstructed SMFs (top) and UVLFs (bottom) for the progenitors of various present-day substructure types, without (solid curves) and with (dashed curves) detectability constraints (Section VIII). The isolated galaxy suite is shown in the left column; paired galaxies in the right column. The proto-MW/LG (black), which includes the progenitors of the present-day main host(s), is compared with reconstructions from the progenitors of satellite galaxies (red), stellar streams (blue), phase-mixed galaxies (green), and field galaxies (yellow).
V.1 Reconstructed differential SMFs and UVLFs for each present-day classification
Figure 4 compares the differential SMFs (top panel) and UVLFs (bottom panel) of the proto-MW/LG (black) at $z=7$ with the ones constructed from the progenitors of present-day satellite galaxies (red), stellar streams (blue), phase-mixed galaxies (green), and field galaxies (yellow). The solid lines show reconstructions from the full set of progenitors and the dashed lines show reconstructions obtained using a simple model for the detectability of present-day substructures with the LSST (section VIII). We show results for the isolated suite and the paired suite in the left and right panels, respectively. The comparison between the two panels illustrates how well the reconstructed SMFs/UVLFs from each classification compare with the proto-MW/LG SMF/UVLF, when the near-far technique is applied to each of the MW and the LG alone. The curves show the median across all of the simulations in the respective galaxy suite. To highlight differences between classifications, the SMFs/UVLFs are computed in the all progenitors scenario rather than the fossil record scenario.